OpenStudy (anonymous):

If V= {p(x) e P3(R); p(-1)=p(1)=0}. Show that B = { 1 - x^2, x - x^3} is a base of V

OpenStudy (anonymous):

If I understand you correctly, you're given that \(V\) the vector space of polynomials \(p(x)\) from the set of degree-(at most)3 polynomials? I'm not exactly up to date on the notation, but I think that's what \(P_3(\mathbb{R})\) means. Anyway, to establish that the set of vectors \(B\) forms a basis, you have to show that any vector in the space \(V\) can be written as a linear combination of the vectors in \(B\). Additionally, the vectors in \(B\) must be linearly independent. First, show they're linearly independent: For all \(x\), can you write \[c_1(1-x^2)+c_2(x-x^3)=0\text{ for }c_1,c_2\neq0?\] Dividing through by \((1-x)(1-x)\) might a good first step: \[c_1+c_2x=0\] For the set to span \(V\), you need to show that any polynomial in the space with roots \(\pm1\) can be written as a linear combination of the vectors in \(B\).